OpenStudy (anonymous):

It takes 3 seconds using a power of 35 W to lift a resting 7 kg bowling ball from the ground and place it on a shelf. Ignoring friction, what is the height to which the bowling ball was lifted? A. 735 m B. 15 m C. 1.5m D. 1.7m What formula would i use?

OpenStudy (anonymous):

\[W = \frac{J}{ s }=\frac{ Nm }{ s }\] solving for meter we rearrange the equation as such\[m =\frac{ Ws }{ N }\]

OpenStudy (anonymous):

W is watts, s is seconds, N is newtons

OpenStudy (anonymous):

Oh and \[N = Mg\] Newtons in this equation is equal to the weight which is the force of gravity. This equals mass times the acceleration due to gravity

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

your welcome.

OpenStudy (anonymous):

you're :)